# LLMs as Hackers: Autonomous Linux Privilege Escalation Attacks

### v4

@@ 2.2.2. Automated “Hacking”
Enumeration scripts such as linux-smart-enumeration [^3] https://github.com/diego-treitos/linux-smart-enumeration/tree/master, linPEAS [^4] https://github.com/carlospolop/PEASS-ng/tree/master/linPEAS or linenum.sh are rule-based. If paths are hard-coded, even simple obfuscation, e.g., installing tools in different locations or running services on untypical ports, can avoid detection. In addition, those tools lack situational awareness, i.e., they are not able to automatically integrate information within found documents, e.g., analyzing a stored email for saved passwords in it.
3. Methodology
We detail wintermute, our prototype to rapidly evaluate privilege-escalation capabilities of LLMs.
We are using a benchmark suite (Happe and Cito, 2024) that provides these security guarantees, and its test scenarios are detailed in Table 1.
3.1. Benchmark Workflow
@@
Our prototype allows for fully automated evaluation of an LLM’s privilege-escalation capabilities. To achieve this, we instantiate new Linux virtual machines (VMs) for each benchmark and use them as a target for the LLM tested. Each of the generated VMs is secure except for the single vulnerability injected into it. The virtual machines are subsequently used as targets for the configured LLM and privilege attacks are performed (detailed in Section 3.2). After root has been achieved or a predefined number of rounds have been reached, the attacks are stopped and the respective VM destroyed. We keep the log information according to Section 3.3 for later analysis.
@@
We make use of VMs as they allow for full control of the target environment. In addition, they provide a good security boundary both between the different test VMs as well as between the benchmark host and the test VMs. As each test-run creates and destroys new VMs, we can ensure that the used VMs are both secure and not tainted by prior runs.
3.2. Prototype: Wintermute
@@
Figure 1. Relationship between prompts and stored data.
You are a low-privilege user \${conn.username} with password \${conn.password} on a \${system} system and your goal is to become the root user by abusing commands and misconfigurations on the tested system. \${capabilities} \${history} \${state} \${guidance} Give your command. Do not add any explanation or an initial \$’. Figure 2. Prompt used to query next command. Your current list of known facts relevant for privilege escalation: \${facts} You executed the command \${cmd}’ yielding the following result: ~~~ bash \${resp} ~~~ Create a new list of facts that unifies the current list of facts as well as new information that you were able to retrieve from the executed command. Try to keep the list of new facts as concise as possible. Figure 3. Prompt used to update the state/worldview.
@@
Wintermute is a Python program that supervises and controls the privilege-escalation attempts. It connects to the target VM through SSH as well as to the used LLMs through an OpenAI-compatible HTTP API. It is also responsible for collecting and storing log information for subsequent analysis.
@@
In wintermute, all decision-making is performed by LLMs through prompting. The main prompt, next-cmd, detailed in Figure 3, includes information currently known by the agent and asks for the next command to be executed on the target virtual machine. The prompts are implemented as Python Mako templates where context information and the current agents worldview are injected through text variables. The interactions between the prompts and the stored data are shown in Figure 1.
@@
The variable capabilities contains a list of available capabilities, i.e., the actions that an LLM can request to interact with the target system. Our prototype provides two actions: execute_command and test_credentials. When using the former, the LLM provides the command to be executed on the target system as parameters, e.g., “execute_cmd ls”. With the latter, the LLM can test the validity of credentials on the target system, e.g., “test_credentials root password”.
@@
The agent’s world view is represented through the variables history, state and guidance. Each of these optional variables can be configured through wintermute. History contains the output of all prior executed commands, i.e., resembles the shell history as seen within an terminal emulator. It is automatically truncated to fit the used LLM’s context size.
@@
The variable state is created using the optional second prompt (update-state) shown in Figure 3. Initially, the state is an empty list. After each executed system command, the LLM is shown the current state, the executed command and its output. It is then tasked with generating a new state, i.e., an updated state list containing information that it has gathered about the target system. If state is used instead of history, the used context size should thus be minimized at the cost of performing an additional LLM prompt as now both next-cmd and update-state are called for each round. This should be suitable for long-running agents where otherwise the size of the accumulated history would incur substantial costs.
Finally, guidance can be used to prime the LLM towards a designated area or vulnerability class. If used, this guidance is statically configured by wintermute for each test scenario.
3.3. Collected Metrics

### v5

@@ 2.2 Penetration Testing
Enumeration scripts such as linux-smart-enumeration [^15] Also often called lse.sh, https://github.com/diego-treitos/linux-smart-enumeration/tree/master, linPEAS [^16] https://github.com/carlospolop/PEASS-ng/tree/master/linPEAS or linenum.sh are rule-based. If paths are hard-coded, even simple obfuscation techniques, e.g., installing tools in different locations or running services on atypical ports, can avoid vulnerability detection. In addition, those tools lack situational awareness, i.e., they are not able to automatically integrate information within found documents, e.g., analyzing a stored email for saved passwords therein.
Benchmarks and Testbeds. In addition to the lack of established Linux Privilege Escalation standards, there is also a lack of Linux Privilege Escalation benchmarks. We assume that one of the reasons is the competitive nature of security testing: as soon as a benchmark is established, tools can optimize for their test-cases, and thus invalidate the benchmark. Security professionals often call this a Red Queen’s Race (Harang and Ducau, 2018).
Due to the sensitive, unpredictable, and potentially destructive nature of security testing experiments, the security and safety of the testbed is of high importance. The commands executed within the test environment must not interact with any non-test system nor network. To achieve this, the test scenarios should be hosted within virtual machines upon a virtual network that is not publicly reachable. This safety requirement, in addition to their ever-changing nature, makes the reuse of online CTF exercises problematic.
2.3 Offensive usage of LLMs for “hacking”
Table 1: Survey Papers used as seed for our Literature Research
Name Authors A Comprehensive Overview of Large Language Models (LLMs) for Cyber Defences Hassanin and Moustafa (2024) A survey on large language model (LLM) security and privacy Yao et al. (2024) From LLMs to LLM-based Agents for Software Engineering Jin et al. (2024) Generative AI in Cyber Security of Cyber Physical Systems Mavikumbure et al. (2024) Large Language Models for Cyber Security: A Systematic Literature Review Xu et al. (2024a) Large Language Models in Cybersecurity: State-of-the-Art Motlagh et al. (2024) Large language models in information security research: A january 2024 survey Dube (2024) LLMs for Intelligent Software Testing: A Comparative Study Boukhlif et al. (2024) Review of Generative AI Methods in Cybersecurity Yigit et al. (2024) When LLMs Meet Cybersecurity: A Systematic Literature Review Zhang et al. (2024)
The potential of LLMs is seen by ethical hackers and blackhats. Gupta et al. (2023) identify multiple areas of interest for using LLMs including phishing/social engineering, pen-testing and the generation of malicious code/binaries (e.g., payloads, ransomware or malware).
Usage by blackhat hackers. Recent darknet monitoring (Gatlan, 2023) indicates that Black-Hats are already offering paid-for LLMs: One suspected threat actor is offering WormGPT (Mascellino, 2023) and FraudGPT: while the former focuses upon social engineering, the latter aids writing malicious code, malware, payloads. The same threat actor is currently preparing DarkBert (Montalbano, 2023) which is supposedly based on the identically named DarkBERT (Jin et al., 2023), a LLM that was designed to combat cybercrime. Other darknet vendors also offer similar products: XXXGPT is advertised for malicious code creation, WolfGPT is advertised for social engineering (Dutta, 2023). Please note that all those products are offered within the darknet behind paywalls, so their claims cannot be independently verified. To the best of our knowledge, there is currently no darknet-offered LLM-aided autonomous penetration testing tool. But, as other areas indicate, their surfacing is just a matter of time.
To gather the state-of-the-art on using LLMs for offensive security, we analyzed recent survey papers highlighted in Table 1 and identified English papers that were using LLMs to perform offensive security in a penetration-testing context. We analyzed citations to pin-down the initial papers that utilized LLMs for offensive security research, resulting in both wintermute and pentestGPT. wintermute (Happe and Cito, 2023a) utilizes a single LLM-driven control loop [^17] Contemporary reseearch into usage of the ReAct pattern uses a similar control patterns (Dagan et al., 2023). to autonomously perform Linux Privilege-Escalation Attacks against a vulnerable CTF-style Linux Virtual Machine containing multiple vulnerabilities. No explicit Task Planner was utilized. To allow for long-running hacking engagements, an optional Summarization module can be used to compact state and context. LLMs measured were GPT-4 and GPT-3.5-turbo of which only GPT-4 allows for successful escalation attacks. All source code and prompts were released on github as open source.
pentestGPT by Deng et al. (2023) utilizes multiple LLM modules181818(Wang et al., 2023) contemporarly name this pattern plan-and-execute. More recent papers prefer the term LLM Agent to LLM module. to solve CTF-style challenges. In contrast to wintermute, they include a human-in-the-loop which executes the tasks given by the LLM. Human operators are allowed limited agency to correct LLM-given commands and analyze the respective outputs of hacking tools, thus making this a non-autonomous system. In another contrast to wintermute, pentestGPT splits exploitation command generation into two distinct modules: a Reasoning Module and a Generation Module. The former introduces a Pentest Task Tree to provide high-level guidance for the penetration test. The reasoning module uses this data to select the next avenue of attack which is then forwarded to the generation module to generate one or more exploitation commands. This split was introduced to allow pentestGPT better usage of long-term memory, to “not fall into rabbit holes”, and allow the LLM to better investigate multiple attack vectors instead of repeating already tried attacks over and over again.
pentestGPT uses a subset of 10 Hack the Box online virtual machines as benchmark and testbed. They utilize GPT-3.5, GPT-4 and Google Bard during their benchmarks. Due to the paper’s age, GPT-3.5 had a context limit of 8k while GPT-4 had a context size limit of 32k imposed upon it. These limits have subsequently been lifted by OpenAI. Out of the 10 virtual machines, GPT-4 was able to successfully exploit 6 ($60\%$) [^19] The paper is not fully consistent. We’re using the summary results from its Table 5, where 6 machines were mentioned as exploited while the data itself suggests that only 5 machines were exploited. machines while GPT-3.5 was only able to exploit 2 ($20\%$) machines. As GPT-3.5 was only able to solve easy challenges, the authors recommend using GPT-4 over GPT-3.5 as the latter “leads to failed tests in simple tasks”. In addition, the paper investigates the usage of vector databases to increase the available context-size but found neglectable impact thus indirectly validating the impact of the high-level Planner control loop to provide long-term planning.
Subsequently mentioned papers cite wintermute, pentestGPT or both of them.
PenHeal by Huang and Zhu (2024) autonomously detects vulnerabilities as well as provides mitigation for the found vulnerabilities. As “only” the vulnerability detection part is relevant to our research, we will focus our review upon it. They utilize a pentestGPT-style, autonomous, high-level architecture with both a Planner and Executor LLM module. They improve upon pentestGPT by adding external knowledge by an additional Instructor module, and by encouraging the LLM to explore multiple diverse attack paths through Counterfactual Prompting (He et al., 2022) [^20] In addition, they mention the use of Roleplay Prompting (Kong et al., 2023) which papers such as AutoAttacker find beneficial as it might circumvent LLMs’ security and policy checks.. The Instructor module is inserted between the Planner and Executor module, and adds penetration testing knowledge before the Executor generates the to-be-executed exploitation commands. It performs Retrieval Augmented Generation (Lewis et al., 2020) based upon two hacking books, Penetration-Testing: A Hands-On Introduction to Hacking (Weidman, 2014) and Metasploit Penetration Testing Cookbook (Singh et al., 2018).
As testbed and benchmark, they use a publicly available CTF-style vulnerable Linux virtual machine, called Metaploitable2. This virtual machine is commonly used by Metasploit training, walk-throughs as well as within one of the books used as base for the Instructor module. The paper does not elaborate upon how memory effects (inclusion of the concrete test vulnerabilities within the LLM or RAG training data) and thus over-training were prevented. They evaluate both GPT-3.5 and GPT-4, of which only GPT-4 is able to achieve successful exploitation.
Through ablation studies they verify the positive impact of both Counter-factional Reasoning and providing external pen-test knowledge. By disabling single features, their respective success rates drop to a third when compared to enabled features.
While their prototype’s source code is not provided nor linked within their paper, detailed examples of their used prompts, their testbed and training documents are.
In “LLM Agents can Autonomously Exploit One-day Vulnerabilities [^21] Please note, that “One-day” is not a “0-day” but denotes previously known vulnerabilities with a known exploit.Fang et al. (2024a) use LLMs to autonomously attack and exploit vulnerabilities. They use a benchmark of 15 test-cases, each of them containing a well-known public vulnerably described through a CVE. They utilize 10 different LLM-configurations for testing, including GPT-3.5, GPT-4, multiple LLama2 models, Mistral/Mixtral, and Hermes-based models. When given detailed information, including exploit code, of the to-be-expected vulnerability, GPT-4 was able to successfully exploit $87\%$ of the test cases while neither of the other models were able to achieve any successful exploitation. This indicates that given a vulnerability description, GPT-4 is able to abuse the vulnerability. When the agent was tasked without prior knowledge of the abusable vulnerability, success rates dropped to $7\%$. While the utilized vulnerabilities within the benchmark set are described, neither source code nor prompts are described within the paper or are released as source code, reducing the reproducibility of the experiment. The author states security reasons for this.
In a parallel paper, “LLM Agents can Autonomously Hack Websites” (Fang et al., 2024b), the authors apply a similar approach towards hacking web applications. To improve the LLM’s knowledge, they add five documents focused upon general web hacking, SQLi, XSS and SSRF. They omit a list of specific included documents due to security reasons. They subsequently test the same 10 LLMs against a synthetic test bed consisting of 15 web-centric security vulnerabilities. Of the tested LLMs, only the GPT-4 based agent was able to exploit $42.7\%$ of the vulnerabilities contained within the benchmark. Through a ablation study they detect that removing either the provided background knowledge, or omitting an unspecified “detailed system instruction prompt” roughly cuts the success rate in half. Neither benchmark, source code, prompts nor the “detailed system instruction” prompt are detailed in the paper (or released as open source) thus reducing the academic reproducibility. One unanswered question is, why in this web-based setting GPT-4 was able to hack $44.7\%$ of test-cases without prior knowledge while in the original paper only a $7\%$ success-rate was achieved.
The limited disclosure of data prevents comparison to prior work. Due to the mentioned techniques and included graphs within the paper, we assume that they implement an architecture similar to wintermute.
In their latest paper “Teams of LLM Agents can Exploit Zero-Day Vulnerabilities” (Fang et al., 2024c) they focus upon hacking web-applications without prior knowledge of the underlying vulnerabilities. They implement a pentestGPT-like Planner pattern for high-level task organisation and use a separate LLM agent (Team Manager) to select one of six task-specific LLM agents. The task-specific agents either specialized on a attack vector class (XSS, SQLi, CSRF, SSTI), were a web vulnerability scanner (ZAP), or an unspecified “generic” web hacking agent. Similarly to their prior work, they added “5–6” undisclosed documents with background-information to the respective task-specific agents. In addition, they mention that they used the same prompt template “but modified them for each vulnerability” (no further description given).
They created a benchmark consisting of 15 well-known web vulnerabilities. To prevent memory effects, only vulnerabilities which were released after the LLMs cut-off date were included. As long as the vulnerabilities were not disclosed on hacking web sites before, this prevents the LLM from “knowing” these vulnerabilities beforehand. As a baseline they used the GPT-4 based agent from their initial paper. This agent has prior knowledge of the to-be-exploited vulnerability and thus should be better able to exploit it. Based upon their prior research, they limited they LLM-selection to GPT-4. The GPT-4 based baseline was able to exploit ${}30\%$ of the vulnerabilities while the new multi-LLM agent was able to abuse ${}19\%$ [^22] As the exact percentages were reported through bars in diagrams, there is a uncertainty due to guessing the exact value from the graphs.. Through an ablation study, they detected that removing the task-specific agents reduces the success rate to roughly a quarter. Keeping the task-specific undisclosed documents from the task-specific agents also reduced the success-rate to a quarter. No ablation study of reducing the task-based agents while keeping the task-specific documents as background information was performed. Neither the prompts, the task-specific documents nor the source code of the prototype were detailed due to security reasons.
Overall, while the three papers are highly relevant to our research topic, their impact is highly reduced by not disclosing essential parts of their solutions and benchmarks.
AutoAttacker (Xu et al., 2024b) uses LLMs to perform autonomous post-breach actions, making its scope similar but broader to wintermute’s scope. In contrast to wintermute, its objectives contain tasks that are not privilege escalation attacks; in addition its custom benchmark contains both Windows and Linux machines.
AutoAttacker implements a pentestGPT-like Planner Architecture and extends it by two additional LLM modules/agents: an Experience Manager and a Navigator module. The Experience Manager utilizes RAG to store information about prior executed attacks. It’s input is the Planner-derived generated hacking tasks; its output are $k$ potential hacking action examples. The orginal Planner-derived hacking task as well as the generated hacking action examples are subsequently presented to the Navigator module which selects the hacking action to execute [^23] The paper does not state if the Experience Manager is reinitialized for each objective or benchmark run, or if it “learns” the hacking objectives and tasks over time over multiple objectives and/or benchmark runs..
Of the benchmarks 14 implemented objectives, at least one objective depends upon executing actions on multiple machines (“Pass-the-Hash”) making this benchmark the first benchmark in which the hacking agent has to connect to multiple targets while solving a single objective. The objectives themselves range from very simple (“File Writing”) to complex (“Pass-the-Hash”). Each task is provided with an “abstract” objective. Given the example in the paper’s appendix, even the abstract objective contains a detailed high-level description of the task at hand, i.e., “…use the Administrator hash to implement a pass-the-hash attack by psexec module…”. The abstract objectives are more detailed than the prompts used by pentestGPT or wintermute. It is more detailed than the objectives given by pentestGPT’s Planner Module to the Execution module, warranting the question why AutoAttacker’s Planner-Module is necessary.
They evaluated GPT-3.5, GPT-4 and LLama2 (both 7b and 70b variants). GPT-4 was able to successfully exploit all given objectives ($100\%$) while GPT-3.5 was only able to solve 3 out of the 14 objectives ($21\%$). Llama2 models were not able to solve a single objective. Through a minimal single-case ablation study the authors show that, e.g., the “Ransomware” objective is only solvable while employing the Experience Manager.
The paper was the only paper that mentioned being detected by policy/ethics filter and details using Roleplaying Prompts (Kong et al., 2023) as countermeasure. While the paper does not contain nor link the benchmark’s or prototype’s source code, the paper details used benchmark cases as well as the utilized prompts.
2.4 Differentiation
Table 2: Related Work
Name Autonomous Code/Prompts available? Reproducibility
Deng et al. (2023) no yes human interaction needed
Fang et al. (2024a) yes no no sources/prompts
Fang et al. (2024b) yes no no sources/prompts
Fang et al. (2024c) yes no no sources/prompts
Xu et al. (2024b) yes only example prompts no sources
Huang and Zhu (2024) yes only example prompts no sources
wintermute (this paper) yes yes
Our work is a continuation of our initial (minimal) prototype of wintermute (Happe and Cito, 2023a), where we now focus on fully-autonomous penetration testing.
This work focuses upon minimizing costly and inefficient LLM module calls. Our prototype’s baseline architecture consists of a control loop utilizing a single LLM invocation, while our more advanced architecture (“state-compaction”) utilizes two LLM callouts within the control-loop. In contrast, most pentestGPT-derived architectures utilize at least a Planner, Executor and Summarizer module while more advanced architectures include Instructor, Experience Manager and Navigator modules.
We focus on Linux Privilege Escalation Attacks as those can be safely be executed within local virtual machines while offering diverse attack paths. In contrast to PenHeal, we focus upon finding a single exploitable vulnerability thus putting us nearer to Penetration Testing than Vulnerability Assessment. While Fang et al. describe these CTF-like challenges as “toy problems”, recent research (Section 2.2) shows that Security Practitioners spent the majority of their time solving these “toy problems”. Coming from a penetration-testing background ourselves, we find the challenge of aiding the vast majority of security professionals worth our time.
We release all our artifacts as open source on github, containing the prototype source-code, benchmark, prompts as well as example run logs. Our benchmark is able to operate offline, thus is heeding our safety requirements. Through publishing it on github, we have already seen usage by third parties. In addition, by creating custom benchmark test-cases from scratch, we prevent contamination and memory effects, e.g., prevent inclusion of benchmark data or walkthroughs within our tested model’s training data.
Exising work has highlighted multiple conflicting opinions regarding the efficacy of utilizing larger context-sizes for history as well the impact of including background hacking information. We will analyze these within our paper through empirical research.
Finally, by focusing on the efficacy of generating and executing hacking commands (Execute LLM model in pentestGPT-derived architectures), we allow future high-level Planners to make better educated choices about the abstraction level of hacking tasks that can be delegated to efficient single-control loop Execution LLM modules. We integrated a “guidance” mechanism into our prototype to be able to emulate both human feedback as well as guided prompting utilized by autonomous LLM Planner agents.
3 Methodology
We detail wintermute, our prototype to rapidly evaluate privilege-escalation capabilities of LLMs. We are using a Linux Privilege Escalation benchmark suite detailed in Section 4 that provides these security guarantees, and its test scenarios are detailed in Table 3. In addition to our specialized Linux Priv-Esc research area, other benchmarks are not publicly available for local testing or do not provide fine-grained vulnerabilities to better analyze LLM behavior.
3.1 Benchmark Workflow
@@
Our prototype allows for fully automated evaluation of an LLM’s privilege-escalation capabilities. To achieve this, we instantiate new Linux virtual machines (VMs) for each new benchmark run. Each of the generated VMs is secure except for the single vulnerability injected into it. The virtual machines are subsequently used as targets for the configured LLM and privilege attacks are performed as detailed in Section 3.2. After root-level access has been achieved, or a predefined number of rounds has been reached, the attacks are stopped and the respective VM destroyed. We keep the log information according to Section 3.3 for later analysis.
@@
We make use of VMs as they allow for full control of the target environment. In addition, they provide a good security boundary between the different test VMs as well as between the benchmark host and the test VMs. As each test-run creates and destroys new VMs, we can ensure that the used VMs are both secure and not tainted by prior runs.
3.2 Autonomous Hacking Agent
@@
Figure 1: Relationship between prompts and stored data. Blue boxes are LLM-callouts, orange boxes denote python modules.
Figure 2: Relationship between prompts and stored data.
You are a low-privilege user \${conn.username} with password
\${conn.password} on a \${system} system and your goal is to
become the root user by abusing commands and misconfigurations
on the tested system.
\${capabilities}
\${history}
\${state}
\${guidance}
Give your command. Do not add any explanation or an initial \$’.
Figure 3: Prompt used to query next command.
Your current list of known facts relevant for privilege escalation:
\${facts}
You executed the command \${cmd}’ yielding the following result:
~~~ bash
\${resp}
Create a new list of facts that unifies the current list of facts
as well as new information that you were able to retrieve from the
executed command. Try to keep the list of new facts as concise as
possible.
Figure 4: Prompt used to update the state/worldview.
@@
hackingBuddyGPT is a Python program that supervises and controls the privilege-escalation attempts. It connects to the target VM through SSH as well as to the used LLMs through an OpenAI-compatible HTTP API. It is also responsible for collecting and storing log information for subsequent analysis.
@@
All decision-making is performed by LLMs through prompting. The main prompt, next-cmd, detailed in Figure 3, includes information currently known by the agent and asks for the next command to be executed on the target machine. The prompts are implemented as Python Mako templates where context information and the current agents worldview are injected through text variables. The interactions between the prompts and the stored data are shown in Figure 1.
@@
Figure 1 also includes the variable capabilities contains a list of available capabilities, i.e., the actions that an LLM can request to interact with the target system. Our prototype provides two actions: execute_command and test_credentials. When using the former, the LLM provides the command to be executed on the target system as parameters, e.g., “execute_cmd ls”. With the latter, the LLM can test the validity of credentials on the target system, e.g., “test_credentials root password”.
@@
The agent’s world view is represented through the variables history, state and guidance. Each of these optional variables can be configured through configuration options. History contains the output of all prior executed commands, i.e., resembles the shell history as seen within an terminal emulator. It is automatically truncated to fit the used LLM’s context size.
@@
The variable state is created using the optional second prompt “update-state” shown in Figure 4. Initially, the state is an empty list. After each executed system command, the LLM is shown the current state, the executed command and its output. It is then tasked with generating a new state, i.e., an updated state list containing information that it has gathered about the target system. If state is used instead of history, the used context size should thus be minimized at the cost of performing an additional LLM prompt as now both next-cmd and update-state are called for each round. This should be suitable for long-running agents where otherwise the size of the accumulated history would incur substantial costs. This implements a Reflection or Iterated Amplification pattern.
Finally, guidance can be used to prime the LLM towards a designated area or vulnerability class. This was implemented to emulate the high-level guidance typically provided to the Executor module in Planner architectures.
3.3 Collected Metrics